If a borrower can afford to repay a home loan by making 300 monthly payments of $986.33, how much can be borrowed at 18% per annum compounded monthly?
1
Expert's answer
2020-04-22T14:17:53-0400

"X=S\frac{i}{(1-(1+i)^{-n})}"


"986.33=S\frac{0.015}{1-(1+0.015)^{-300}}"


"986.33=S\frac{0.015}{1-(1+0.015)^{-300}}=S\times0.015=65 032.02"


S=65 032.02